# Efficient Reasoning Models: A Survey

### v1

@@ 2.2 Why We Need Efficient Reasoning
Efficiency is a valuable research direction across many fields, and in the context of reasoning, we highlight key motivations for pursuing efficient reasoning (see Figure 3). Reasoning models often generate excessively long reasoning chains to solve reasoning tasks, even for simple samples (see Appendix A.1 for a concrete example), and typically rely on larger model sizes to achieve stronger reasoning performance. Additionally, some strategies, such as Best-of-N and self-consistency (Wang et al., 2022), further scale the decoding process to enhance reasoning performance. These lead to substantial computational and memory demands. Moreover, overly long reasoning paths can accumulate errors and negatively impact final accuracy (Wu et al., 2025c; Yang et al., 2025c).
On the other hand, efficient reasoning is also essential in real-world applications such as embodied AI (Duan et al., 2022), agent systems (Wang et al., 2024a), and real-time platforms (e.g., autonomous driving (Cui et al., 2024)). In these scenarios, efficiency enables agents to process sensory inputs in real time, make swift and accurate decisions, and interact seamlessly with dynamic environments. Additionally, unnecessarily lengthy reasoning may increase safety risks (Kuo et al., 2025; Li et al., 2025c), posing unpredictable threats. These challenges collectively highlight the limitations of current reasoning models, underscoring the necessity of improving reasoning efficiency.
@@ 3.1 Make Long CoT Short
CoT prompting has significantly improved the reasoning capabilities of LLMs by explicitly generating intermediate reasoning steps. However, generating overly long CoTs can lead to negative issues (see Section 2.2). Recent works have explored various approaches to mitigate these drawbacks by shortening CoT length without compromising reasoning performance. Among them, RL with length penalty is widely used for encouraging concise and effective reasoning paths (see Section 3.1.1). Another line of work explores SFT with variable-length CoT data to improve reasoning efficiency, as discussed in Section 3.1.2. In addition, prompt-driven techniques improve reasoning efficiency by utilizing prompts, with further details available in Section 3.1.3. Finally, we explore latent reasoning, which performs the reasoning process in latent space and drastically reduces CoT length, with details provided in Section 3.1.4. Additionally, Table 1 provides an overview of these methods. Furthermore, we present the performance of various efficient reasoning methods on the challenging AIME dataset in Table 5.
@@
[…]
@@
3.1.1 Reinforcement Learning Helps Efficiency Improvement
Incorporating explicit chain length penalties into RL is a natural strategy for shortening reasoning chains (Team et al., 2025; Li et al., 2025a; Arora & Zanette, 2025). L1 (Aggarwal & Welleck, 2025) takes this further by introducing designated length-constraint instructions into the training data. O1-Pruner (Luo et al., 2025a) develops a specialized reward design by utilizing length and accuracy from a reference model as baselines, explicitly rewarding shorter reasoning paths and higher accuracy to ensure efficiency without sacrificing performance. DAST (Shen et al., 2025b) aims to achieve a balanced CoT (i.e., dynamically adjusting computational resources by allocating more reasoning steps to more challenging questions and fewer to simpler ones). Specifically, it proposes a Token Length Budget (TLB), defined as a weighted sum of the mean token count in accurate answers and a predefined upper bound on generation length to quantify problem difficulty, penalizing excessively verbose reasoning for simple questions while encouraging comprehensive reasoning for complex ones. THINKPRUNE (Hou et al., 2025) designs a length-aware reward function that only provides a reward if the correct answer is generated within a specified token budget. The model is trained using the Group Relative Policy Optimization (GRPO) algorithm with progressively tightened length constraints. Additionally, Think When You Need (Yang et al., 2025b) utilizes pairwise comparisons to generate rewards based on the relative length and accuracy of reasoning, guiding models to produce concise yet accurate solutions.
Fine-tuning LLMs using variable-length CoT data effectively teaches models to solve reasoning tasks with shorter paths. Following a clear fine-tuning pipeline, we organize the discussion of this line of research into two stages: (1) how variable-length CoT data is constructed and (2) which SFT approach (i.e., standard or progressive) is adopted. For each work, we explicitly address these two questions to facilitate comparison and analysis.
@@ 3.1.2 Supervised Fine-Tuning with Variable-Length CoT Data Helps Efficiency Improvement
Another line of works adopts progressive fine-tuning strategies (Liu et al., 2024b; Ma et al., 2025). LLM-Skip (Liu et al., 2024b) iteratively encourages the model to generate shorter reasoning paths and then merges the generated shorter paths into the training set for subsequent fine-tuning rounds, gradually reducing chain length. CoT-Valve (Ma et al., 2025) supports both standard SFT and two progressive strategies: CoT-Valve++ and CoT-Valve+P. CoT-Valve++ introduces a normalized path-length factor $\beta$, which is smaller for longer paths. During training, the model parameters are dynamically adjusted along a direction scaled by $\beta$, allowing the model to adapt to reasoning paths of varying lengths and learn finer-grained length control. CoT-Valve+P, on the other hand, progressively trains the model on samples sorted from long to short chains, guiding it to shorten the chain length over successive fine-tuning stages.
@@ 3.1.3 Prompt-Driven Efficiency Enhancement in Reasoning
Prompt-driven techniques have emerged as a flexible approach to improving reasoning efficiency. We categorize related works into two directions: (1) prompt-guided reasoning, which leverages well-designed prompt to guide reasoning models toward more effective reasoning paths and (2) prompt-based routing, which utilizes prompt-level attributes (e.g., complexity) to adaptively select appropriate computational paths (i.e., route easy questions to lightweight models and hard ones to powerful large models).
@@
This line of work focuses on enhancing reasoning efficiency through prompt engineering, encouraging the model to produce concise reasoning paths while preserving performance. Concise CoT (Renze & Guven, 2024) shows that simply adding “Be concise” to the prompt can shorten reasoning chains. Break the Chain (Ding et al., 2024) leverages carefully crafted instructions (e.g., “rapidly evaluate and use the most effective reasoning shortcut”) to trigger the model’s ability to exploit shortcuts and skip unnecessary steps. TALE-EP (Han et al., 2024) employs an LLM-based estimator to predict the minimal token budget required for each question, which is then incorporated into the prompt to guide efficient reasoning. CoD (Xu et al., 2025b) develops the instruction “Think step by step, but only keep a minimum draft for each thinking step, with 5 words at most.”, which significantly reduces token usage under few-shot settings without compromising accuracy. However, its performance degrades in zero-shot settings and on small language models. MARP (Chen et al., 2024a) boosts per-step information density and reduces step count under a fixed reasoning boundary, achieving high efficiency gains through prompt design, and can be further combined with PoT for better computation-reasoning separation. Token-Complexity (Lee et al., 2025) presents token complexity to measure the minimal tokens needed for correct reasoning and derives the theoretical compression limit of CoT chains. Through prompt variations (e.g., “use 10 words or less” or “remove all punctuation”), they explore the trade-off between performance and efficiency and show that current methods still fall far from the optimal bound, leaving room for improvement. Additionally, these methods can effectively construct variable-length CoT data, thereby supporting the approaches introduced in Section 3.1.2.
@@
This line of work improves reasoning efficiency by leveraging prompt attributes (e.g., question complexity) to design routing strategies that dynamically adapt inference models. Claude 3.7 Sonnet (Anthropic., 2025) offers two response modes (e.g., quick answers or step-by-step thinking), allocating more compute to complex reasoning tasks. Although the implementation details remain undisclosed, it is the first hybrid reasoning model and a foundation for subsequent methods.
@@ 3.1.4 Reasoning in Latent Space
Unlike explicit CoT reasoning, latent reasoning (Deng et al., 2023) performs the reasoning process entirely in the model’s latent space, skipping the generation of explicit intermediate steps. Latent reasoning brings two key benefits: it allows for more human-like thinking by modeling complex ideas beyond language and improves efficiency by reducing the need for explicit reasoning chains. This section first examines how models transition from explicit to implicit reasoning. Then, we explore how reasoning is represented in latent space (e.g., discrete tokens or continuous tokens).
@@
Another line of work explores using an auxiliary model to generate latent reasoning tokens directly from the input. CCoT (Cheng & Van Durme, 2024) trains a lightweight CCOT module (a LoRA (Hu et al., 2022)) to produce compressed latent reasoning tokens directly from input, which are then fed into a decoding module to generate concise answers, while HCoT (Liu et al., 2024c) adopts a similar pipeline but places greater emphasis on semantic alignment during compression. SoftCoT (Xu et al., 2025c) adopts a similar strategy by training a lightweight assistant model to produce implicit representations conditioned on the input. Furthermore, Reasoning with Latent Thoughts (Saunshi et al., 2025) demonstrated that looping a transformer multiple times could emulate a deeper model and naturally induce latent thoughts, effectively capturing iterative reasoning without tokenized steps. RELAY (Yu et al., 2025a) follows this idea by aligning each iteration of a looped transformer (Giannou et al., 2023) with explicit CoT steps. The trained looped model is then leveraged to produce high-quality CoT chains to train stronger autoregressive models on long reasoning tasks.
@@
A common choice for latent space representation is to use continuous tokens (Zhang et al., 2025a; Shen et al., 2025c; Cheng & Van Durme, 2024; Xu et al., 2025c; Hao et al., 2024; Liu et al., 2024c), which naturally align with the internal computation of neural networks. Coconut (Hao et al., 2024) models reasoning in the hidden space by feeding the final-layer hidden states back into the model without decoding explicit CoT tokens, enabling more continuous and efficient reasoning. This approach unlocks advantages that explicit CoT cannot offer, such as backtracking and parallel decoding. Inspired by Coconut, Heima (Shen et al., 2025a) introduces thinking tokens into multimodal large language models (MLLMs) to replace explicit reasoning steps, enabling reasoning in latent space.
@@ 3.2 Build Small Language Model with Strong Reasoning Ability
An alternative approach to improving reasoning efficiency is to empower small language models (SLMs) with strong reasoning capabilities. Due to their lower memory and computational requirements, SLMs are inherently more efficient and easier to deploy in real-world applications. Model compression (Han et al., 2016; Frantar et al., 2023; Li et al., 2023b) naturally aligns with this goal, as it enables small or compressed models to retain or gain reasoning abilities. A natural starting point is to transfer reasoning capabilities from larger models via distillation (see Section 3.2.1). We further explore other model compression techniques, including pruning and quantization, which aim to compress models without severely compromising reasoning performance in the Section 3.2.2. Beyond traditional model compression techniques, RL offers another promising direction, enhancing reasoning capabilities under limited resources through carefully designed training strategies, as discussed in Section 3.2.3. Additionally, a summary of these methods is presented in Table 2.
Apart from model compression and RL, some studies explore the reasoning ability of small language models from alternative perspectives. For example, Liu et al. (2025d) shows that small language models can match or even surpass the reasoning performance of much larger LLMs with carefully designed TTS strategies. However, these TTS strategies are not universally applicable, and the optimal approach depends heavily on factors such as the model architecture, reward model design, and problem complexity. While small language models have shown impressive reasoning capabilities, they often face challenges in instruction following and self-reflection, indicating that further adaptation or fine-tuning may still be required to better align the model’s behavior with human intent.
@@
[…]
@@ 3.2.1 Distillation Transfers Reasoning Ability to Small Language Model
CoT-KD (Magister et al., 2022) first demonstrated that distillation can transfer reasoning ability from LLMs to small language models. However, due to limited capacity, small language models struggle to learn complex reasoning (Li et al., 2025d), motivating the development of more advanced strategies. Based on the optimization target, existing methods can be grouped into two directions: (1) data-focused, which improves the quality or composition of training data and (2) model-focused, which concentrates on the distilled model itself or its generation strategy.
@@
MD (Li et al., 2023a) adopts mix distillation by combining data generated with different prompting strategies (CoT and PoT) as training data, and Mix (Li et al., 2025d) applies a similar strategy using a mix of long and short CoT samples. CD (Feng et al., 2024) enhances training diversity by mixing original data with counterfactual samples derived from it, while NAT (Li et al., 2024a) leverages negative data. DLCoT (Luo et al., 2025c) improves training data quality by segmenting and simplifying long reasoning paths. SCORE (Zhang et al., 2024) enables self-correction by allowing the model to generate, identify, and refine its reasoning, using the corrected outputs for further distillation. Distill2-to-1 (Yu et al., 2024) only retrains (input, answer) pairs as training data. The above methods rely on standard SFT, but some adopt progressive SFT. FDD (Zhu et al., 2024b) progressively adjusts data difficulty based on the small language model’s performance on LLM-generated data, while SKIntern (Liao et al., 2025b) proposes a progressive process that removes symbolic knowledge and examples step by step, encouraging the model to internalize reasoning ability.
@@
PRR (Zhao et al., 2024) distills two separate models: a probing model for retrieving relevant knowledge and a reasoning model for generating answers based on the question and retrieved content. Thinking slow, fast (Paliotta et al., 2025) explores distilling reasoning ability from transformer-based models into Mamba or Mamba-Transformer architectures to reduce inference cost to lower costs. Additionally, ATM (Chen et al., 2024b) designs an adaptive mechanism that enables the student model to dynamically choose between pre-thinking (i.e., thinking before answering) and post-thinking (i.e., answering before thinking) based on question complexity.

### v2

@@ 2.5 Why We Need Efficient Reasoning
Efficiency is a valuable research direction across many fields, and in the context of reasoning, we highlight key motivations for pursuing efficient reasoning (see Figure 3). Reasoning models often generate excessively long reasoning chains to solve reasoning tasks, even for simple samples, and typically rely on larger model sizes to achieve stronger reasoning performance. For example, answering “What is the answer of 1 plus 2?” requires 619 tokens from DeepSeek R1-685B (see Appendix A.2 for details). To further illustrate the overhead, we evaluated four versions of DeepSeek R1 on the AIME 24 dataset and observed consistently huge token counts: 15513 for 1.5B, 12377 for 7B, 10854 for 14B, and 10024 for 32B. Additionally, some strategies, such as Best-of-N and self-consistency, further scale the decoding process to enhance reasoning performance. These lead to substantial computational and memory demands. Moreover, overly long reasoning paths can accumulate errors and negatively impact final accuracy (Wu et al., 2025d; Yang et al., 2025c).
On the other hand, efficient reasoning is also essential in real-world applications such as embodied AI (Duan et al., 2022), agent systems (Wang et al., 2024a), and real-time platforms (e.g., autonomous driving (Cui et al., 2024)). In these scenarios, efficiency enables agents to process sensory inputs in real time, make swift and accurate decisions, and interact seamlessly with dynamic environments. Additionally, unnecessarily lengthy reasoning may increase safety risks (Kuo et al., 2025; Li et al., 2025d), posing unpredictable threats. These challenges collectively highlight the limitations of current reasoning models, underscoring the necessity of improving reasoning efficiency.
Table 1: Performance of efficient reasoning methods on the AIME 24 dataset. $\dagger$ denotes the result of the original model, averaged over 5 independent runs.
Category Type Methods Acc. / #Tokens Base Model
Original Model - Baseline 70.67% / 10024 DeepSeek-R1-32B
Shorter RL DAST 53.30% / 6337 DeepSeek-R1-Distill-Qwen-7B
Shorter SFT CoT-Valve 43.30% / 4630 QwQ-32B-Preview
Shorter SFT TOPS 46.00% / 6427 Qwen2.5-32B
Smaller KD Mix 10.00% / - Qwen2.5-3B
Smaller KD DLCoT 53.30% / 18825 Qwen2.5-14B
Smaller RL Open-RS 46.70% / - DeepSeek-R1-Distill-Qwen-1.5B
Smaller RL DeepSacleR 43.10% / - DeepSeek-R1-Distill-Qwen-1.5B
Faster Efficient self-consistency RPC 9.50% / - InternLM-2-MATH-Plus 7B
Faster Efficient sampling $\phi$-Decoding 16.67% / - LLaMA3.1-8B-I
@@ 3 Efficient Reasoning
In the following, we introduce efficient reasoning methods based on three key categories: shortening long chains of thought, as discussed in Section 3.1; developing small language models with strong reasoning capabilities, details of which can be found in Section 3.2; and improving decoding efficiency, which is elaborated in Section 3.3. We present the performance of various efficient reasoning methods on the challenging AIME 24 dataset in Table 1 and further provide a latency-based summary of representative methods across categories on the GSM8K dataset in Table 5.
@@ 3.1 Make Long CoT Short
Recent works have explored various approaches to improve reasoning efficiency by shortening CoT length without compromising reasoning performance. Among them, RL with length penalty is widely used for encouraging concise and effective reasoning paths (see Section 3.1.1). Another line of work explores SFT with variable-length CoT data to improve reasoning efficiency, as discussed in Section 3.1.2. In addition, prompt-driven techniques improve reasoning efficiency by utilizing prompts, with further details available in Section 3.1.3. Finally, we explore latent reasoning, which performs the reasoning process in latent space and drastically reduces CoT length, with details provided in Section 3.1.4. Additionally, Table 2 provides an overview of these methods, showing that most RL-based methods utilize Full FT, while many SFT-based methods adopt Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA (Hu et al., 2022) to reduce cost. This trend suggests that RL-based methods require more extensive parameter updates, making lightweight adaptation less effective; for latent reasoning, Full FT remains dominant, and these methods often yield higher speedups, indicating that implicit representations enable more effective compression and offer a higher upper bound compared to explicit reasoning chains.
3.1.1 Reinforcement Learning Helps Efficiency Improvement
Incorporating explicit chain length penalty into RL is a natural strategy for shortening reasoning chains (Team et al., 2025; Li et al., 2025a; Arora & Zanette, 2025). L1 (Aggarwal & Welleck, 2025) takes this further by introducing designated length-constraint instructions into the training data. O1-Pruner (Luo et al., 2025a) develops a specialized reward design by utilizing length and accuracy from a reference model as baselines, explicitly rewarding shorter reasoning paths and higher accuracy to ensure efficiency without sacrificing performance. DAST (Shen et al., 2025b) aims to achieve a balanced CoT (i.e., dynamically adjusting computational resources by allocating more reasoning steps to more challenging questions and fewer to simpler ones). Specifically, it proposes a Token Length Budget (TLB), defined as a weighted sum of the mean token count in accurate answers and a predefined upper bound on generation length to quantify problem difficulty, penalizing excessively verbose reasoning for simple questions while encouraging comprehensive reasoning for complex ones. THINKPRUNE (Hou et al., 2025) designs a length-aware reward function that only provides a reward if the correct answer is generated within a specified token budget. The model is trained using the Group Relative Policy Optimization (GRPO) algorithm with progressively tightened length constraints. Additionally, Think When You Need (Yang et al., 2025b) utilizes pairwise comparisons to generate rewards based on the relative length and accuracy of reasoning, guiding models to produce concise yet accurate solutions.
@@ 3.1.2 Supervised Fine-Tuning with Variable-Length CoT Data Helps Efficiency Improvement
Latent reasoning Implicit-KD SFT (Full FT) GSM8K: 20.00% / - GPT-2 small 8.2 $\times$ (L.)
@@
Following a clear fine-tuning pipeline, we organize the discussion of this line of research into two stages: (1) how variable-length CoT data is constructed and (2) which SFT approach (i.e., standard or progressive) is adopted. For each work, we explicitly address these two questions to facilitate comparison and analysis.
@@
Another line of work adopts progressive fine-tuning strategies (Liu et al., 2024b; Ma et al., 2025). LLM-Skip (Liu et al., 2024b) iteratively encourages the model to generate shorter reasoning paths and then merges the generated shorter paths into the training set for subsequent fine-tuning rounds, gradually reducing chain length. CoT-Valve (Ma et al., 2025) supports both standard SFT and two progressive strategies: CoT-Valve++ and CoT-Valve+P. CoT-Valve++ introduces a normalized path-length factor $\beta$, which is smaller for longer paths. During training, the model parameters are dynamically adjusted along a direction scaled by $\beta$, allowing the model to adapt to reasoning paths of varying lengths and learn finer-grained length control. CoT-Valve+P, on the other hand, progressively trains the model on samples sorted from long to short chains, guiding it to shorten the chain length over successive fine-tuning stages.
@@ 3.1.3 Prompt-Driven Efficiency Enhancement in Reasoning
We categorize prompt-driven works into two directions: (1) prompt-guided reasoning, which leverages well-designed prompts to guide reasoning models toward more effective reasoning paths and (2) prompt-based routing, which utilizes prompt-level attributes (e.g., complexity) to adaptively select appropriate computational paths (e.g., route easy questions to lightweight models and hard ones to powerful large models).
@@
Concise CoT (Renze & Guven, 2024) shows that simply adding “Be concise” to the prompt can shorten reasoning chains. Break the Chain (Ding et al., 2024) leverages carefully crafted instructions (e.g., “rapidly evaluate and use the most effective reasoning shortcut”) to trigger the model’s ability to exploit shortcuts and skip unnecessary steps. TALE-EP (Han et al., 2024) employs an LLM-based estimator to predict the minimal token budget required for each question, which is then incorporated into the prompt to guide efficient reasoning. CoD (Xu et al., 2025c) develops the instruction “Think step by step, but only keep a minimum draft for each thinking step, with 5 words at most.”, which significantly reduces token usage under few-shot settings without compromising accuracy. However, its performance degrades in zero-shot settings and on small language models. MARP (Chen et al., 2024a) boosts per-step information density and reduces step count under a fixed reasoning boundary, achieving high efficiency gains through prompt design, and can be further combined with PoT for better computation-reasoning separation. Token-Complexity (Lee et al., 2025) presents token complexity to measure the minimal tokens needed for correct reasoning and derives the theoretical compression limit of CoT chains. Through prompt variations (e.g., “use 10 words or less” or “remove all punctuation”), they explore the trade-off between performance and efficiency and show that current methods still fall far from the optimal bound, leaving room for improvement. Additionally, these methods can effectively construct variable-length CoT data, thereby supporting the approaches introduced in Section 3.1.2.
@@
Claude 3.7 Sonnet (Anthropic., 2025) offers two response modes (e.g., quick answers or step-by-step thinking), allocating more compute to complex reasoning tasks. Although the implementation details remain undisclosed, it is the first hybrid reasoning model and a foundation for subsequent methods.
@@ 3.1.4 Reasoning in Latent Space
Unlike explicit CoT reasoning, latent reasoning (Deng et al., 2023; Tan et al., 2025) performs the reasoning process in latent space, skipping the generation of explicit intermediate steps. Latent reasoning brings two key benefits: it allows for more human-like thinking by modeling complex ideas beyond language, and improves efficiency by reducing the need for explicit reasoning chains. This section first examines how models transition from explicit to implicit reasoning. Then, we explore how reasoning is represented in latent space.
@@
Another line of work explores using an auxiliary model to generate latent reasoning tokens directly from the input. CCoT (Cheng & Van Durme, 2024) trains a lightweight CCOT module (a LoRA (Hu et al., 2022)) to produce compressed latent reasoning tokens directly from input, which are then fed into a decoding module to generate concise answers, while HCoT (Liu et al., 2024c) adopts a similar pipeline but places greater emphasis on semantic alignment during compression. SoftCoT (Xu et al., 2025d) adopts a similar strategy by training a lightweight assistant model to produce implicit representations conditioned on the input. Furthermore, Reasoning with Latent Thoughts (Saunshi et al., 2025) demonstrated that looping a transformer multiple times could emulate a deeper model and naturally induce latent thoughts, effectively capturing iterative reasoning without tokenized steps. RELAY (Yu et al., 2025a) follows this idea by aligning each iteration of a looped transformer (Giannou et al., 2023) with explicit CoT steps. The trained looped model is then leveraged to produce high-quality CoT chains to train stronger autoregressive models on long reasoning tasks.
@@
A common choice for latent space representation is to use continuous tokens (Zhang et al., 2025a; Shen et al., 2025c; Cheng & Van Durme, 2024; Xu et al., 2025d; Hao et al., 2024; Liu et al., 2024c), which naturally align with the internal computation of neural networks. Coconut (Hao et al., 2024) models reasoning in the hidden space by feeding the final-layer hidden states back into the model without decoding explicit CoT tokens, enabling more continuous and efficient reasoning. This approach unlocks advantages that explicit CoT cannot offer, such as backtracking and parallel decoding. Inspired by Coconut, Heima (Shen et al., 2025a) introduces thinking tokens into multimodal large language models (MLLMs) to replace explicit reasoning steps, enabling reasoning in the latent space.
@@ 3.2 Build Small Language Model with Strong Reasoning Ability
Compared to compressing reasoning chains, an alternative approach to improving reasoning efficiency is to empower small language models (SLMs) with strong reasoning capabilities. Due to their lower memory and computational requirements, SLMs are inherently more efficient and easier to deploy in real-world applications. Model compression (Han et al., 2016; Frantar et al., 2023b; Li et al., 2023b) naturally aligns with this goal, as it enables small or compressed models to retain or gain reasoning abilities. A natural starting point is to transfer reasoning capabilities from larger models via distillation (see Section 3.2.1). We further explore other model compression techniques, including pruning and quantization, which aim to compress models without severely compromising reasoning performance in Section 3.2.2. Beyond traditional model compression techniques, RL offers another promising direction, enhancing reasoning capabilities under limited resources through carefully designed training strategies, as discussed in Section 3.2.3. Additionally, a summary of these methods is presented in Table 3, indicating that most distillation approaches still rely on Full FT, with a few adopting PEFT techniques. Notably, methods that progressively incorporate refined or synthesized data (e.g., FDD and SKIntern) tend to achieve greater performance improvements.
Apart from model compression and RL, some studies explore the reasoning ability of small language models from alternative perspectives. For example, Liu et al. (2025d) shows that small language models can match or even surpass the reasoning performance of much larger LLMs with carefully designed TTS strategies. However, the effectiveness of TTS strategies varies with model architecture, reward design, and task complexity. While small language models show potential in reasoning, their limitations in instruction following and self-reflection highlight the need for further adaptation to align with human intent.
@@
RL Open-RS GRPO (Full FT) $\text{Blended}_{1}$ AIME: 46.70% ($\uparrow$ 17.80%) DeepSeek-R1-Distill-Qwen-1.5B
@@ 3.2.1 Distillation Transfers Reasoning Ability to Small Language Model
CoT-KD (Magister et al., 2022) first demonstrated that distillation can transfer reasoning ability from LLMs to small language models. However, due to limited capacity, small language models struggle to learn complex reasoning (Li et al., 2025e), motivating the development of more advanced strategies. Based on the optimization target, existing methods can be grouped into two directions: (1) data-focused, which improves the quality or composition of training data, and (2) model-focused, which concentrates on the distilled model itself or its generation strategy.
@@
MD (Li et al., 2023a) adopts mix distillation by combining data generated with different prompting strategies (CoT and PoT) as training data, and Mix (Li et al., 2025e) applies a similar strategy using a mix of long and short CoT samples. CD (Feng et al., 2024c) enhances training diversity by mixing original data with counterfactual samples derived from it, while NAT (Li et al., 2024a) leverages negative data. DLCoT (Luo et al., 2025c) improves training data quality by segmenting and simplifying long reasoning paths. SCORE (Zhang et al., 2024) enables self-correction by allowing the model to generate, identify, and refine its reasoning, using the corrected outputs for further distillation. Distill2-to-1 (Yu et al., 2024) only retrains (input, answer) pairs as training data. The above methods rely on standard SFT, but some adopt progressive SFT. FDD (Zhu et al., 2024b) progressively adjusts data difficulty based on the small language model’s performance on LLM-generated data, while SKIntern (Liao et al., 2025b) proposes a progressive process that removes symbolic knowledge and examples step by step, encouraging the model to internalize reasoning ability.
@@
PRR (Zhao et al., 2024) distills two separate models: a probing model for retrieving relevant knowledge and a reasoning model for generating answers based on the question and retrieved content. Thinking slow, fast (Paliotta et al., 2025) explores distilling reasoning ability from transformer-based models into Mamba or Mamba-Transformer architectures to reduce inference cost. Similarly, M1 (Wang et al., 2025b) builds on Mamba (Gu & Dao, 2024) to develop a hybrid linear RNN reasoning model that alleviates latency and memory overhead from long reasoning chains, further enhanced through RL after distillation. Additionally, works such as NSA (Yuan et al., 2025) and MoBA (Lu et al., 2025), which focus on lightweight architectures for general efficiency, can also be extended to improve reasoning efficiency. Additionally, ATM (Chen et al., 2024b) designs an adaptive mechanism that enables the student model to dynamically choose between pre-thinking (i.e., thinking before answering) and post-thinking (i.e., answering before thinking) based on question complexity.